Reports 1-1 of 1 Clear search Modify search
MIF (General)
tomotada.akutsu - 21:17 Tuesday 03 August 2021 (17748) Print this report
Shut the green beam on the POP table for investigating FNC

Follow-up of 17733

As entitled. To investigate what caused the fringe amplitude change, we started with shutting one of the Michelson arms on the POP table by putting a beam dump in front of the fiber output.

Images attached to this report
Comments to this report:
takahiro.yamamoto - 11:16 Wednesday 04 August 2021 (17755) Print this report
As expected, though, the laser power at the output of Prometheus was not fluctuated.

During last night, a beam dump shut out also the return light to PSL room.
So a PD at PSL room received just only the output light of Prometheus.

Akutsu-san put the beam dump around -17h in Fig.1.
Then, the fluctuation in readout value of the PD for X disappeared.

Fringe amplitude may be affected by fiber coupling at PSL.
By the way, I found K1:ALS-?_FIB_SERVO_.* was set up farther downstream than I had expected.
A figure in Mirror map became slightly old.
JGW-D1808971 is newer than Mirror map. But I heard it was not also the latest layout.
Images attached to this comment
tomotada.akutsu - 21:05 Wednesday 04 August 2021 (17769) Print this report

So the fluctuation around 0 during shutting the green beam retuning from POP would just represent a fluctuation of P/S ratio in the reference arm in the PSL room? Anyway, surprisingly the fluctiation amplitude ~1000 counts is nearly a half of the "healthy" amount of of the fringe (~ 2000 counts)

Search Help